Operations Report – December 04th

Crew 216 Operations Report 04-Dec-19
SOL: 10
Name of person filing report: Mike Lawson
Non-nominal systems: NA
Generator: run
Hours run: 11.5
From what time last night: 1916
To what time this morning: 0651
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 82 %
Diesel Reading – 34
Station Propane Reading – 52
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1464901
Water (static tank) (gal): 214
Static to Loft Pump used – yes
Water in Green Hab (gal): 251
Water in Science Dome (gal): 0
Toilet tank emptied: no
Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: no
Hours: 116
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging:
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: no
Hours: 121
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: Opportunity to be moved off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments: EVA to Phobos Peak. HAB local EVA for ISS comms test. No contact with ISS.
Summary of internet: No serious issues.
Summary of suits and radios: Suit 4 airflow is weak, possibly a battery charging issue.
Summary of Hab operations: Routine maintenance and housekeeping tasks. Hot water line vibration is extreme when pump engages. Recommend securing it to HAB wall or risk fitting failure.
Summary of GreenHab operations: routine watering.
Summary of Science Dome operations: Test ingots of plaster of paris poured for future testing.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved.
Summary of any observatory issues: NA
Summary of health and safety issues: NA
Questions, concerns and requests to Mission Support: NA

Crew 216 Journalist Report 04Dec2019

[title Journalist Report– December 04th]

Crew 216 Journalist Report 04-DEC-19
Sol 10
Author: Evgenia Alexandrova

Mornings begin really early in the Hab. Usually around 5am one can see the glimpses of light under the doors of certain crew members. Jetlag doesn’t give up on me so at this time I usually read. This morning’s reading brought this insight to me. There are two main things that are really important in life. First: quality of our relationships with other people. Second: whether or not we will have enough courage to admit who we really are and make our dreams come true. At MDRS we practice both. While we play board games and when we climb a rock on an EVA. When we cook diner and when we try to contact the ISS. When we water the plants in the GreenHab and when we work on gypsum in the Science Dome. Even when we just wait in silence the end of depressurization in the Airblock.

And traditionally a new poem for the audience:

Walking on the sand to roam

We found an ancient blackened bone

Further still were polished stones

By wind and moving water born

We reached the mighty Phobos peak

And all around saw majesty

The ground was smooth and rough between

We marveled at the things we’d seen

Crew 216 Astronomy Report 04Dec2019

[title Astronomy Report – December 04th]

Crew 216 Astronomy Report 04-DEC-19
Crew Astronomer: Michael Ho

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: none
Objects to be Imaged this Evening: M1, NGC 7635, NGC 281
Images submitted with this report: Crab_191108
Problems Encountered: nil

MUSK OBSERVATORY
Solar Features Observed: None.
Images submitted with this report: nil
Problems Encountered: Clouds

"You cannot change the course of the wind, but you can adjust your sails."

Crew 216 Greenhab Report 04Dec2019

[title Greenhab Report – December 04th]

Crew 216 GreenHab Report 04-DEC-19
Crew GreenHab Officer: Michael Ho

Environmental control: Door opened for ventilation at 12.15 pm, closed at 5.00 pm.
Average temperature: 22C
Hours of supplemental light: Automatic system
Daily water usage for crops: 20L
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 10.10 am; 4.30 pm.

SEASONAL PLANTING HISTORY:
(Plants introduced since 03/11/2019 till 22/11/2019 please refer to Crew 215 GreenHab Report of 22/11/2019.)
Crew 216:
26/11/2019:
1 avocado seed in small black pot
24/11/2019:
1 avocado seed in small black pot
23/11/2019:
seeded Zucchini in 2 large black pots
seeded Chilli Pepper ‘Shishito’ in 2 2-gallon black pots
seeded Sweet Greenfield Pepper ‘California Wonder’ in 2 2-gallon black pots
seeded Swiss Chard ‘Flamingo Pink’ in 4 2-gallon black pots
transplanted three self-seeded tomato plants in large blue tub transplanted to 2-gallon pots

OTHER NOTES:
1st sprouts:
Zucchini – 2nd of 2 pots!

Harvest: None
Support/supplies requested:nil.

Crew 216 EVA Report 04Dec2019

[title EVA Report – December 04th]

Crew 216 EVA Report 04-Dec-2019
EVA # 14
Author: Kennedy
Purpose of EVA: To capture film footage in support of Evgenia’s and Ho’s projects.
Start time for EVA: 0830
End time for EVA: 1030
Narrative:Left the Hab on foot with Kennedy, Alexandrova, Levesque, and Ho. We traveled due east towards Phobos peak. In transit we discovered a large dinosaur bone protruding from the side of a cliff. GPS coordinates taken and coordinated with Shannon. We then continued to the southern base of Phobos ridge. Michael stayed at the bottom to run his drone footage. We climbed to the top of the ridge with drone footage and pictures taken from the Hab. We reached the peak around 0945. After returning to the base we returned to the Hab following a more northerly route, passing by several capstone formations “mushrooms” passing in site of the marble ritual, before returning to the Hab at 1030. Both Alexandrova and Ho reported excellent footage being taken for their respective projects.
Destination:Phobos Peak
Coordinates (use UTM NAD27 CONUS): 12 S 0519319E 4250181N
EVA Participants: Alexandrova, Kennedy, Ho, Levesque
Road(s) and routes per MDRS Map:
Road(s) and routes per MDRS Map:
1) Walked to from Hab to base of Phobos Peak
2) Ascended Phobos Peak
2) Returned to Hab from Phobos Peak
Mode of travel: Foot
Vehicles used: None

[end]

Crew 216 Research Report 04Dec2019

[title Research Report – December 04th]

Crew 216 Research Report 4-Dec-2019
Sol 10
Submitted by: Whittle

1) Psych Scale (TAMU IRB2019-1042D) – 3rd stage surveys completed, final stage on Saturday.

2) Medical Casts – Created 6 sample blocks with varying composition, 1:2 water to gypsum and 7.5:10 respectively. Sample blocks currently drying. Developed methodology for shear stress testing. Will test tomorrow to record shear stress. More powdered gypsum ready for use with simulated cast using strongest composition.

3) In-Situ Fluorescent Mineral Prospecting – EVA was successful with several samples obtained. Further evaluation will be done in Hab in a controlled dark room and compared against reference materials.

4) MDRS Mapping – Nothing further to report.

5) MDRS Communications – No new developments on this project today. Attempted communications with ISS flyover, no response.

6) Mars terraforming simulation was executed with success in 18 generations. Kennedy’s corporation was ultimately successful in the technical aspects but the whole group was able to enjoy the process.

[end]

Crew 216 Commander Report 04Dec2019

[title Commander Report – December 04th]

Crew 216 Commander Report 04-DEC-2019
Sol: 10
Summary Title: High and Low
Author: Marc Levesque
When you walk away from the Hab, you are always treated to something special. The rocks, soils, shapes, colors, and sun angle form a unique optical combination no matter where you look. The people who chose the MDRS location nailed it for an analog Martian landscape. Walking around in an EVA suit adds to this feeling that you truly are on a different planet. Today’s trip to the summit of Phobos Peak for Evgenia’s filming and Michael’s drone footage once again reinforced how special it is to be here at MDRS and participate as a crew in this environment. After decompressing from this high, we then prepared for our attempt to contact the International Space Station via amateur radio. The ISS orbital schedule had it passing directly overhead at 1315 for 10 minutes. After we set up my vehicle’s mobile radio as a cross-band repeater, we returned to the Hab to attempt the contact. We knew it would be a very long shot, as there were too many variables that had to come together to make it happen. Most crucial of all was that an astronaut had to be listening, which is something they do only on an occasional basis unless a contact is set up months in advance through a formal program. The crew was pretty excited by the prospect, however, and while we were disappointed that no contact was made, all felt the effort worthwhile. From a ham’s perspective, it was a great technical exercise.

Crew 216 EVA Report 04Dec2019

[title EVA Report – December 04th]

Crew 216 EVA Report 04-Dec-2019
EVA # 15
Author: Levesque
Purpose of EVA: To set up a cross-band repeater in the Crew Commander’s vehicle for an attempt to contact the International Space Station via amateur radio.
Start time for EVA: 1230
End time for EVA: 1345
Narrative: Walked to vehicle to set up radio in cross-band mode. Returned to Hab for contact attempt, and then returned to vehicle to shut down radio.
Destination: Crew Commander’s vehicle next to Outpost.
Coordinates (use UTM NAD27 CONUS): 12 S 0518364 4249273.
EVA Participants: Levesque and Lawson
Road(s) and routes per MDRS Map:
1) Walk from Hab to vehicle to set up the repeater
2) Return to Hab for contact attempt with a handheld radio
3) Return to vehicle to shut off repeater and the vehicle
Mode of travel: Foot

[end]

Crew 216 Sol Summary 04Dec2019

[title Sol Summary – December 04th]

Crew 216 Sol Summary Report 04-DEC-19
Sol: 10
Summary Title: Phobos Peak and International Space Station
Author: Marc Levesque
Mission Status: Active
Sol Activity Summary: The morning’s EVA to Phobos Peak captured some great film footage for Evgenia’s project and allowed Michael an opportunity to do the same for his drone project. The landscape around MDRS is different no matter which direction you go, and the view from Phobos was spectacular. The afternoon EVA was short, less than 100 meters, to set up the radio in the Crew Commander’s vehicle for an attempt to conduct an amateur radio contact with the International Space Station. Alas, no one was listening, but it was certainly worth a shot. As usual, GreenHab and Astronomy activities were ongoing.
Look Ahead Plan: Starting to wrap up projects with no EVAs.
Anomalies in work: N/A.
Weather: Clear and cold.
Crew Physical Status: Healthy
EVAs: To Phobos Peak and Crew Commander’s vehicle
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, Astronomy Report, and Photos.
Support Requested: None

[end]